For a solution of acetic acid (HC2H3O2) to be called vinegar it must contain 5% acetic acid by mass. If the vinegar is made up of only of acetic acid, density of the solution = 1.006g/ml. Find mol fraction of solute and solvent, M, m of the solution.



1
Expert's answer
2021-04-14T06:13:44-0400

Solution:

Vinegar is an aqueous solution of acetic acid.

acetic acid - HC2H3O2 = C2H4O2

The molar mass of C2H4O2 is 60.052 g/mol.

The molar mass of H2O is 18.015 g/mol.


Let's assume that the volume of the solution is 1 L (1000 mL).

Thus,

Mass of 1 L C2H4O2 solution = Solution volume × Density = 1000 mL × 1.006 g/mL = 1006 g

Mass of C2H4O2 = (%C2H4O2 × Mass of solution) / 100% = (5% × 1006 g) / 100% = 50.3 g C2H4O2

Mass of solvent (H2O) = Mass of solution - Mass of C2H4O2 = 1006 g - 50.3 g = 955.7 g H2O

Moles of C2H4O2 = Mass of C2H4O2 / Molar mass of C2H4O2 = 50.3 g / 60.052 g mol-1 = 0.8376 mol

Moles of H2O = Mass of H2O / Molar mass of H2O = 955.7 g / 18.015 g mol-1 = 53.0502 mol


Mole Fraction (χ):

Mole fraction (χ) is the ratio of moles of one substance in a mixture to the total number of moles of all substances.

Total number of moles = Moles of C2H4O2 + Moles of H2O = 0.8376 mol + 53.0502 mol = 53.8878 mol

χsolute = χC2H4O2 = Moles of C2H4O2 / Total number of moles = 0.8376 mol / 53.8878 mol = 0.016

χsolute = 0.016

χsolvent = χH2O = Moles of H2O / Total number of moles = 53.0502 mol / 53.8878 mol = 0.984

χsolvent = 0.984


Molarity (M):

Molarity or molar concentration is the number of moles of solute per liter of solution.

Molarity of solution = Moles of C2H4O2 / Volume of solution

Molarity of solution = 0.8376 mol / 1 L = 0.8376 mol/L = 0.8376 M = 0.838 M

Molarity of solution = 0.838 M


Molality (m):

Molality of a solution is the moles of solute divided by the kilograms of solvent.

Mass of solvent (H2O) = 955.7 g = 0.9557 kg

Molality of solution = Moles of C2H4O2 / Kilograms of solvent

Molality of solution = 0.8376 mol / 0.9557 kg = 0.8764 mol/kg = 0.8764 m = 0.876 m

Molality of solution = 0.876 m


Answers:

χsolute = 0.016

χsolvent = 0.984

Molarity of solution = 0.838 M

Molality of solution = 0.876 m